OpenStudy (anonymous):

factoring Trinomials 2x^2-11x-6

OpenStudy (anonymous):

(2x+1)(x-6)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

ywc

OpenStudy (anonymous):

Is there anyway you can show me the steps so I can get a better picture please.

OpenStudy (anonymous):

steps factoring Trinomials?

OpenStudy (anonymous):

yes please or this equation.

OpenStudy (anonymous):

it hard I wil ltry explain

OpenStudy (anonymous):

Okay thank you.

OpenStudy (anonymous):

2x*x =2x^2 -6=1*-6 or: -1*6

OpenStudy (anonymous):

how would you factor out 20c^2+13c-15 as an example, if you don't mind me asking

OpenStudy (anonymous):

I have tried to do this on a graphing calculator but i am confused on how to do it

OpenStudy (saifoo.khan):

.

OpenStudy (saifoo.khan):

u want the answers right?

OpenStudy (saifoo.khan):

x = 3/5 x = -5/4

OpenStudy (anonymous):

20c^2= 1c*20C 2C*10C 4c*5c -15= -1*15 or -15*1 or 3*-5 or -5 *3

OpenStudy (anonymous):

I just want a better understanding on how to do it, because my professor is confusing me.

OpenStudy (anonymous):

okay thank you everyone I am sure I will be back on. :)

OpenStudy (anonymous):

4c 5 5c - 3 cross and time: 4c*-3= -12c 5c*5 = 25c 25c-12c = 13c (4c+5)(5c-3)

OpenStudy (anonymous):

Thank you, you have been a big help.

OpenStudy (anonymous):

r u understand?

OpenStudy (anonymous):

Yes I think I got it. lol I havn't done algebra in almost 7 years so I am refreshing my mind.

OpenStudy (anonymous):

good